CURRICULUM VITAE of Domokos SZÁSZ

Personal: Born: August 18, 1941, Budapest. Married, 3 children

Position: Research Professor

Mathematical Institute of the Hungarian Academy of Sciences H-1364 Budapest P.O.B. 127 Hungary Phone: (36–1) 117-7175 or (36–1) 117-3050 (Secr.)

Record: 1964:  Diplom in Mathematics, Eötvös University, Budapest
1964-68: Assistant Professor at the Probability Department

of Eötvös University, Budapest

1967:  Ph.D., Eötvös University, Budapest
1969:  Grünwald Prize
1971:  Candidate’s degree (with distinction), Lomonosov University, Moscow
1971—: Present affiliation
1981:  Degree of the Doctor of Mathematical Sciences
1984:  1984 Prize of the Hungarian Academy of of Sciences
1990:  Corresponding Member of the Hungarian Academy of Sciences
1993-95: Director of the Institute
1995:  Ordinary Member of the Hungarian Academy of Sciences
1995:  Szele Prize
Memberships: Ed. Boards of Journal of Statistical Physics (1989-1991),

Ergodic Theory and Dynamical Systems (1995-)

of Studia Sci. Math. Hung. (Chief Editor, 1993-96) and of Acta Math. Acad. Sci Hung., Intn’l Sci. Adv. Committee of Schrödinger Institute, Vienna Bernoulli Society, ISI, Bolyai János Mathematical Society, IAMP.

Research topics: stochastic processes, dynamical processes, nonequilibrium

statistical physics.

Invited papers (last years):

1. Memorial Conference on the 150th Birthday of L. Boltzmann, February 1994, Vienna 2. Dynamics Days International Conference, June 1994, Budapest

3. Symposium on Classical and Quantum Billiards, July 1994, Ascona, Switzerland 4. Workshop on Dynamical Systems, May 1995, Trieste, Italy 5. Symposium on Ergodic Theory and Dynamical Systems, July 1995, Warsaw

6. Conference on Dynamical Systems, Oberwolfach, July 1995 and 1997

7. Probability and Physics, Workshop, August 1995, Renkum, Netherlands

8. Mark Kac Lectures, Amsterdam, Spring 1996

9. International Symposium on Dynamical Systems, Rio de Janeiro, August 1997

Academic invitations:

1. IHES, Bures-sur-Yvette, (Visiting Member) 2 months, 1983

2. Università di Roma, Dipartimento di Matematica, 6 weeks, 1987 (CNR Visiting Scholar)

3. Rutgers University, Department of Mathematics, 6 weeks, 1988

4. Princeton University, Department of Mathematics, (Visiting Senior Research Mathematician) and Institute for Advanced Study, School of Mathematics (Visiting Member) 2 months, 1989

5. University of Arizona, Tucson, Department of Mathematics, 4 weeks, 1989

6. Institute for Advanced Study, School of Mathematics (Member), 1990 Fall Semester,

7. IHES, Bures-sur-Yvette, (Visiting Member) 1 month, 1993

8. Instituto de Matematica Pura et Aplicada, Rio de Janeiro, 1 month, April 1995

9. Erwin Schrödinger Insitute, Vienna, 2 months, 1996 Fall

Visiting professorships:

1. Dartmouth College, USA (Visiting Associate Professor), 6 months, 1977

2. Goethe Universität, Frankfurt (Visiting Full Professor) 6 months , 1977-78

3. Dartmouth College, USA (Visiting Full Professor), 6 months, 1985

4. Princeton University, USA (Visiting Full Professor), 1990-91.

Seminars: dynamical systems: Berkeley, Caltech, CUNY, Maryland, Moscow, Princeton,

Tucson, ETH (Zürich), Paris VII., Hebrew University, PennState Univ., etc. statistical physics: Heidelberg, IHES, Moscow, Princeton, Rome, Rutgers, Vienna, Lausanne, etc. probability theory: Caltech, Göttingen, Harvard-MIT, Moscow, NYU, Paris, Princeton, USC, etc.

Organizing activity:

1. Co-chairman of the OC, First and Second Colloquia on Random Fields, 1979 (Esztergom) and 1984 (Kőszeg)

2. Session Organizer, 7th International Congress on Mathematical Physics, 1983, Boulder, Colorado 3. Member the OC, Semester on Dynamical Systems, 1986, Banach Center, Warsaw 4. Member of Advisory Committee, 9-th (1988, Swansee), 10-th (1991, Leipzig), 11-th

(1994, Paris) International Congresses on Mathematical Physics,

5. Co-chairman of the OC, Symposium on Classical and Quantum Billiards, 1994, Ascona, Switzerland.

6. Member of the Organizing Committee, Workshop on Statistical Physics, August 1995, Budapest, Hungary

7. Vice-chairman of the Scientific Committee, European Congress of Mathematicians, Budapest, 1996

8. Chairman of the Scientific Committee, Semester on Hyperbolic Systems with Singularities, Erwin Schrödinger Institute, Vienna, Fall of 1996